OpenStudy (anonymous):

What are the sum of the measures of the interior angles of a convex hexagon?

OpenStudy (jmark):

Sum of the measures of the interior angles of a convex hexagon we know that sum of the measures of the interior angles of a convex polygon with n side= 180(n-2) here we have convex hexagon then n=6 there for Sum of the measures of the interior angles of a convex hexagon=180(6-2) =180(4) =720 Learn more about Area of a Hexagon at http://goo.gl/0XekRx